# Quantifying Robustness: A Benchmarking Framework for Deep Learning Forecasting in Cyber-Physical Systems

### v2

@@ Abstract
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: XX.XXXX/XX
@@ 1 Introduction
Cyber-Physical Systems (CPS) continuously generate large-scale and complex time series data across industrial domains such as manufacturing, process engineering, and energy distribution, which enable vital tasks including Prognostics and Health Management (PHM) [1]. Deep Learning (DL) has proven effective in modeling complex data, yet despite its predictive capabilities the adoption of DL in industrial CPS remains limited due to concerns over robustness [2]. Robustness, in this context, refers to the model’s ability to maintain reliable performance when encountering erroneous or unforeseen inputs [3]. Regulatory developments, such as the European Union’s AI Act, have further highlighted the importance of robust DL systems, urging systematic methodologies for robustness assessment [4]. However, consensus on practical robustness testing methodologies remains elusive [5]. While formal verification methods and adversarial robustness studies have been prominent in the literature, their practical relevance to industrial scenarios is limited due to unrealistic assumptions [6]. Recent research emphasizes evaluating model robustness against realistic data quality issues and distributional shifts commonly encountered in real-world settings [7, 8, 9, 10]. However, existing studies stop short of combining an operational robustness metric and a large-scale evaluation on real CPS data, leaving practitioners without a practical benchmark.
@@ 2 Related Work
Robustness in DL: Robustness of a DL model is generally described as the ability "to maintain its level of performance under any circumstances" [3], or its capability "to cope with erroneous, noisy, unknown, and adversarial input data" [11]. It is closely related to other concepts such as reliability [12], resilience [13], safety [6] and stability [14]. DL models are known to be highly sensitive to small input perturbations[15]. Consequently, much research has focused on adversarial robustness, evaluating model resilience to these small, intentionally crafted perturbations [16], leading to formal guarantees and metrics to measure this type of robustness [17]. However, adversarial perturbations rarely occur naturally, thus recent research emphasizes measuring robustness to broader, more realistic distributional shifts [18, 19]. Traditional adversarial robustness metrics often fail to sufficiently capture these distributional shifts [20, 21]. To address this limitation, recent studies propose evaluating robustness either by assessing performance across multiple distinct environments [22], by applying broad yet realistic natural distribution shifts (e.g., image blur or style changes) [23], or by specifically investigating the impact of data-quality issues [24, 25].
@@
Robustness for Time Series and CPS: Robustness in time series forecasting is generally assessed by evaluating how model performance deteriorates in response to anomalies, commonly measured by increased forecasting errors [26], or by measuring the effect on the output distribution, e.g. by measuring the Wasserstein distance [27]. In CPS, research of robustness has traditionally focused on formal verification [28] or stability in control engineering [14]. Other CPS-focused robustness research has explored adversarial attacks [7], false-positive predictions under varying grades of handicap [29], and model robustness to added random noise [10]. Recent studies have specifically investigated the effects of data quality issues such as sensor drift, outliers and missing values on simulated CPS data [8, 9].
@@
Prognostics in CPS: In CPS contexts, PHM assesses system health and predicts future system behavior [1]. Many PHM applications rely on DL forecasting models, either by using them to directly predict the remaining useful lifetime (RUL) or by first forecasting health indicators of the system. In general, more accurate forecasts of health indicators are strongly associated with improved RUL prediction accuracy [30, 31]. Commonly used architectures include recurrent neural networks (e.g., LSTM [32], GRU [33]), modular approaches like Recurrent Independent Mechanisms (RIMs) [34], Temporal Convolutional Networks (TCN) [35], attention-based models such as Transformers [36] and Informers [37], and structured state-space models like Mamba [38]. Recent findings indicate that even simple feed-forward neural networks can be competitive [39]. These models represent diverse architectural approaches commonly adopted in CPS forecasting tasks. Despite their widespread application, the robustness of these architectures has primarily been evaluated using synthetic or idealized datasets, potentially masking their real-world vulnerabilities.
@@
In summary, no prior paper combines an operational robustness score and a large-scale evaluation on real CPS data under realistic disturbances. Papers that introduce formal definitions validate them only on synthetic case studies [10, 29] or focus solely on adversarial noise rather than real-world scenarios [7]. Studies that do examine realistic data-quality issues limit their evaluations to simulations and report robustness as simple performance deltas [8, 9]. This paper closes these gaps by proposing a distribution-grounded robustness score and benchmarking nine forecasting models under representative sensor-fault scenarios across six industrial CPS data sets.
@@ 3 Defining Robustness for CPS
($,Y)=((\textbf{x}_{i},\dots,\textbf{x}_{i+t-1}),(\textbf{x}_{i+t},\dots,\textbf{x}_{i+t^{\prime}-1}))\in\textbf{X}.$ X . (1)
@@
where $10<\varepsilon<\!<$ 1 is a small constant added for numerical stability. The relative performance $\mu_{rel}$ measures the effect of a disturbance by comparing the performance on the disturbed sample against the original performance. Note that $\mu_{\text{rel}}$ makes no statement about the inherent performance of the model, it only measures the effect of the disturbances.
@@
[…]
@@ 4 Proposed Framework
To evaluate the proposed definition of robustness, we implemented a framework that simulates disturbances in CPS datasets to quantify the robustness of popular DL model architectures. We focus on a forecasting task as a representative objective, since it requires a deep understanding of the underlying system, which can translate into more precise RUL estimates [30, 31]. The framework is openly accessible, supports custom time series data in CSV or Parquet format and can easily integrate models implemented in Python. An overview of the proposed framework is shown in Figure 1.
@@ 5 Experimental Setup
  • Electricity[40]: This dataset provides hourly electricity consumption records for 321 customers over the period from 2012 to 2014.
  • ETT[37]: The ETT dataset captures measurements from electronic transformers, including load and oil temperature, over a two-year period. It is offered at two temporal resolutions: Hourly aggregated data with 17,420 observations across 7 features (ETTh1) and data recorded at 15-minute intervals, comprising 69,680 observations across 7 features (ETTm1).
  • SKAB[41]: The SKAB dataset contains sensor readings from a water circulation system testbed. Data were collected at a one-second sampling rate over a three-hour period, resulting in 9,405 observations of 8 sensor variables.
  • SWaT[42]: The Secure Water Treatment (SWaT) dataset by iTrust, Centre for Research in Cyber Security, Singapore University of Technology and Design consists of sensor and actuator measurements obtained from a scaled-down version of an industrial water treatment plant. Data were recorded at one-second intervals continuously over seven days, yielding 449,919 observations with 52 features.
  • Water Quality[43]: This dataset monitors water quality in the water supply system serving a state in Germany, with 9 sensor readings taken at one-minute intervals from August to November 2016.
@@
  • DLinear[39]: A linear forecasting model that integrates a time series decomposition strategy to separately model trend and seasonal components.
  • MLP[44]: A Multi-Layer Perceptron that employs a series of linear transformations interleaved with ReLU activation functions [45] and batch normalization [46] to enhance convergence and generalization.
  • TCN[35]: A Temporal Convolutional Network that leverages causal convolutions and dilated convolutions to capture long-range dependencies of temporal data.
  • LSTM[32]: A Long Short-Term Memory network that incorporates gating mechanisms to mitigate the vanishing gradient problem.
  • GRU[33]: A Gated Recurrent Unit network that simplifies the LSTM architecture by combining the forget and input gates into a single update gate.
  • RIMs[34]: A Recurrent Independent Mechanisms model that decomposes the hidden state into multiple independent modules. Each module selectively processes different aspects of the input, thereby enhancing robustness.
  • Transformer[36]: A self-attention based architecture that abandons recurrence in favor of parallelized computations.
  • Informer[37]: An efficient variant of the Transformer specifically designed for long sequence forecasting.
  • Mamba[38]: A SSM that exploits linear time-invariant dynamics to efficiently model sequential data.
@@
All models are trained using a batch size of 64. The fixed learning rate is chosen as part of the hyperparameter tuning. The optimization is carried out using the Adam optimizer [47].
@@
All training is executed on a Kubernetes cluster comprising three nodes, each equipped with 128 CPU cores and 500GB of memory, alongside a total of 4 NVIDIA A30 GPUs. Similar to the approach presented in [48], we orchestrated the execution of the experiments using Kubeflow pipelines.
@@ 6.1 Baseline Forecasting Performance
Fully Connected DLinear [39] 0.2174 $\pm$ 0.1393 0.2587 $\pm$ 0.1809
Fully Connected MLP [44] 0.2284 $\pm$ 0.1523 0.3916 $\pm$ 0.2460
Convolution TCN [35] 0.2470 $\pm$ 0.1550 0.4362 $\pm$ 0.2834
Recurrent LSTM [32] 0.2575 $\pm$ 0.1758 0.5016 $\pm$ 0.3676
Recurrent GRU [33] 0.2393 $\pm$ 0.1679 0.4620 $\pm$ 0.3209
Modular RIMs [34] 0.2383 $\pm$ 0.1655 0.3868 $\pm$ 0.2250
Attention Transf. [36] 0.2200 $\pm$ 0.1575 0.3576 $\pm$ 0.2123
Attention Informer [37] 0.2203 $\pm$ 0.1550 0.3552 $\pm$ 0.2006
SSM Mamba [38] 0.3047 $\pm$ 0.1945 0.7177 $\pm$ 0.7299
@@ 6.2 Robustness Evaluation
Fully Connected DLinear [39] 0.2754 $\pm$ 0.2455
Fully Connected MLP [44] 0.4577 $\pm$ 0.2494
Convolution TCN [35] 0.4498 $\pm$ 0.2814
Recurrent LSTM [32] 0.6411 $\pm$ 0.3412
Recurrent GRU [33] 0.5948 $\pm$ 0.3543
Modular RIMs [34] 0.5886 $\pm$ 0.2856
Attention Transf. [36] 0.4745 $\pm$ 0.2577
Attention Informer [37] 0.4748 $\pm$ 0.2670
SSM Mamba [38] 0.5186 $\pm$ 0.3804
@@ 7 Discussion
In this paper, we have introduced a practical definition of robustness tailored specifically to industrial CPS time series and evaluated this definition through a comprehensive forecasting benchmark involving multiple DL architectures. One insight is the strong performance of relatively simple models such as DLinear, confirming earlier findings [39]. However, despite its excellent predictive accuracy on both validation and test datasets, DLinear was highly vulnerable to disturbances, achieving the lowest robustness score among all tested architectures. This shows that standard evaluation methods might not detect overfitting and fail to quantify robustness, which is further affirmed by empirical research in the deployment of machine learning models [49].

### v3

@@ Abstract
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/ETFA65518.2025.11205527
@@ 1 Introduction
Cyber-Physical Systems (CPS) continuously generate large-scale and complex time series data across industrial domains such as manufacturing, process engineering, and energy distribution, which enable vital tasks including Prognostics and Health Management (PHM) [leePrognosticsHealthManagement2014]. Deep Learning (DL) has proven effective in modeling complex data, yet despite its predictive capabilities the adoption of DL in industrial CPS remains limited due to concerns over robustness [windmannArtificialIntelligenceIndustry2024]. Robustness, in this context, refers to the model’s ability to maintain reliable performance when encountering erroneous or unforeseen inputs [ISOIECTR2021]. Regulatory developments, such as the European Union’s AI Act, have further highlighted the importance of robust DL systems, urging systematic methodologies for robustness assessment [RegulationEU20242024]. However, consensus on practical robustness testing methodologies remains elusive [schuettRiskManagementArtificial2024]. While formal verification methods and adversarial robustness studies have been prominent in the literature, their practical relevance to industrial scenarios is limited due to unrealistic assumptions [perez-cerrolazaArtificialIntelligenceSafetyCritical2023]. Recent research emphasizes evaluating model robustness against realistic data quality issues and distributional shifts commonly encountered in real-world settings [zhouRobustnessTestingData2022, dixMeasuringRobustnessML2023, windmannRobustnessGeneralizationPerformance2023, schmidtAssessingRobustnessDataDriven2025]. However, existing studies stop short of combining an operational robustness metric and a large-scale evaluation on real CPS data, leaving practitioners without a practical benchmark.
@@ 2 Related Work
Robustness in DL: Robustness of a DL model is generally described as the ability "to maintain its level of performance under any circumstances" [ISOIECTR2021], or its capability "to cope with erroneous, noisy, unknown, and adversarial input data" [DINSPEC9200112019]. It is closely related to other concepts such as reliability [billintonReliabilityEvaluationEngineering1992], resilience [moskalenkoResilienceResilientSystems2023], safety [perez-cerrolazaArtificialIntelligenceSafetyCritical2023] and stability [runggerNotionRobustnessCyberPhysical2016]. DL models are known to be highly sensitive to small input perturbations[szegedyIntriguingPropertiesNeural2014]. Consequently, much research has focused on adversarial robustness, evaluating model resilience to these small, intentionally crafted perturbations [bastaniMeasuringNeuralNet2016], leading to formal guarantees and metrics to measure this type of robustness [sinhaCertifyingDistributionalRobustness2018]. However, adversarial perturbations rarely occur naturally, thus recent research emphasizes measuring robustness to broader, more realistic distributional shifts [buzhinskyMetricsMethodsRobustness2021, hendrycksUnsolvedProblemsML2022]. Traditional adversarial robustness metrics often fail to sufficiently capture these distributional shifts [jiangFantasticGeneralizationMeasures2019, dziugaiteSearchRobustMeasures2020]. To address this limitation, recent studies propose evaluating robustness either by assessing performance across multiple distinct environments [yuanOutDistributionGeneralization2022], by applying broad yet realistic natural distribution shifts (e.g., image blur or style changes) [hendrycksManyFacesRobustness2021], or by specifically investigating the impact of data-quality issues [schelterJENGAFrameworkStudy2021, mohammedEffectsDataQuality2024].
@@
Robustness for Time Series and CPS: Robustness in time series forecasting is generally assessed by evaluating how model performance deteriorates in response to anomalies, commonly measured by increased forecasting errors [chengRobustTSFTheoryDesign2024], or by measuring the effect on the output distribution, e.g. by measuring the Wasserstein distance [yoonRobustProbabilisticTime2022]. In CPS, research of robustness has traditionally focused on formal verification [mitschModelPlexVerifiedRuntime2016] or stability in control engineering [runggerNotionRobustnessCyberPhysical2016]. Other CPS-focused robustness research has explored adversarial attacks [zhouRobustnessTestingData2022], false-positive predictions under varying grades of handicap [pfrommerReduceHandicapPerformance2023], and model robustness to added random noise [schmidtAssessingRobustnessDataDriven2025]. Recent studies have specifically investigated the effects of data quality issues such as sensor drift, outliers and missing values on simulated CPS data [dixMeasuringRobustnessML2023, windmannRobustnessGeneralizationPerformance2023].
@@
Prognostics in CPS: In CPS contexts, PHM assesses system health and predicts future system behavior [leePrognosticsHealthManagement2014]. Many PHM applications rely on DL forecasting models, either by using them to directly predict the remaining useful lifetime (RUL) or by first forecasting health indicators of the system. In general, more accurate forecasts of health indicators are strongly associated with improved RUL prediction accuracy [wangHealthIndicatorForecasting2020, sunPredictingFutureCapacity2024]. Commonly used architectures include recurrent neural networks (e.g., LSTM [hochreiterLongShortTermMemory1997], GRU [choLearningPhraseRepresentations2014]), modular approaches like Recurrent Independent Mechanisms (RIMs) [goyalRecurrentIndependentMechanisms2021], Temporal Convolutional Networks (TCN) [leaTemporalConvolutionalNetworks2017], attention-based models such as Transformers [vaswaniAttentionAllYou2017] and Informers [zhouInformerEfficientTransformer2021], and structured state-space models like Mamba [guMambaLinearTimeSequence2024]. Recent findings indicate that even simple feed-forward neural networks can be competitive [zengAreTransformersEffective2023]. These models represent diverse architectural approaches commonly adopted in CPS forecasting tasks. Despite their widespread application, the robustness of these architectures has primarily been evaluated using synthetic or idealized datasets, potentially masking their real-world vulnerabilities.
@@
In summary, no prior paper combines an operational robustness score and a large-scale evaluation on real CPS data under realistic disturbances. Papers that introduce formal definitions validate them only on synthetic case studies [schmidtAssessingRobustnessDataDriven2025, pfrommerReduceHandicapPerformance2023] or focus solely on adversarial noise rather than real-world scenarios [zhouRobustnessTestingData2022]. Studies that do examine realistic data-quality issues limit their evaluations to simulations and report robustness as simple performance deltas [dixMeasuringRobustnessML2023, windmannRobustnessGeneralizationPerformance2023]. This paper closes these gaps by proposing a distribution-grounded robustness score and benchmarking nine forecasting models under representative sensor-fault scenarios across six industrial CPS data sets.
@@ 3 Defining Robustness for CPS
$(X,Y)=((\textbf{x}_{i},\dots,\textbf{x}_{i+t-1}),(\textbf{x}_{i+t},\dots,\textbf{x}_{i+t^{\prime}-1}))\in\textbf{X}.$ (1)
@@
where $0<\varepsilon<\!<1$ is a small constant added for numerical stability. The relative performance $\mu_{rel}$ measures the effect of a disturbance by comparing the performance on the disturbed sample against the original performance. Note that $\mu_{\text{rel}}$ makes no statement about the inherent performance of the model, it only measures the effect of the disturbances.
@@
$R_{d}(f,\mu)=\mathbb{E}\left[\int\limits_{0}^{1}\mu_{\text{rel}}(f,d,s,\mathcal{X},\mathcal{Y})ds\right].$ (6)
@@ 4 Proposed Framework
To evaluate the proposed definition of robustness, we implemented a framework that simulates disturbances in CPS datasets to quantify the robustness of popular DL model architectures. We focus on a forecasting task as a representative objective, since it requires a deep understanding of the underlying system, which can translate into more precise RUL estimates [wangHealthIndicatorForecasting2020, sunPredictingFutureCapacity2024]. The framework is openly accessible, supports custom time series data in CSV or Parquet format and can easily integrate models implemented in Python. An overview of the proposed framework is shown in Figure 1.
@@ 5 Experimental Setup
  • Electricity[wuAutoformerDecompositionTransformers2021]: This dataset provides hourly electricity consumption records for 321 customers over the period from 2012 to 2014.
  • ETT[zhouInformerEfficientTransformer2021]: The ETT dataset captures measurements from electronic transformers, including load and oil temperature, over a two-year period. It is offered at two temporal resolutions: Hourly aggregated data with 17,420 observations across 7 features (ETTh1) and data recorded at 15-minute intervals, comprising 69,680 observations across 7 features (ETTm1).
  • SKAB[katserSKABSkoltechAnomaly2020]: The SKAB dataset contains sensor readings from a water circulation system testbed. Data were collected at a one-second sampling rate over a three-hour period, resulting in 9,405 observations of 8 sensor variables.
  • SWaT[gohDatasetSupportResearch2017]: The Secure Water Treatment (SWaT) dataset by iTrust, Centre for Research in Cyber Security, Singapore University of Technology and Design consists of sensor and actuator measurements obtained from a scaled-down version of an industrial water treatment plant. Data were recorded at one-second intervals continuously over seven days, yielding 449,919 observations with 52 features.
  • Water Quality[rehbachGECCO2018Industrial2018]: This dataset monitors water quality in the water supply system serving a state in Germany, with 9 sensor readings taken at one-minute intervals from August to November 2016.
@@
  • DLinear[zengAreTransformersEffective2023]: A linear forecasting model that integrates a time series decomposition strategy to separately model trend and seasonal components.
  • MLP[rumelhartLearningRepresentationsBackpropagating1986]: A Multi-Layer Perceptron that employs a series of linear transformations interleaved with ReLU activation functions [nairRectifiedLinearUnits2010] and batch normalization [ioffeBatchNormalizationAccelerating2015] to enhance convergence and generalization.
  • TCN[leaTemporalConvolutionalNetworks2017]: A Temporal Convolutional Network that leverages causal convolutions and dilated convolutions to capture long-range dependencies of temporal data.
  • LSTM[hochreiterLongShortTermMemory1997]: A Long Short-Term Memory network that incorporates gating mechanisms to mitigate the vanishing gradient problem.
  • GRU[choLearningPhraseRepresentations2014]: A Gated Recurrent Unit network that simplifies the LSTM architecture by combining the forget and input gates into a single update gate.
  • RIMs[goyalRecurrentIndependentMechanisms2021]: A Recurrent Independent Mechanisms model that decomposes the hidden state into multiple independent modules. Each module selectively processes different aspects of the input, thereby enhancing robustness.
  • Transformer[vaswaniAttentionAllYou2017]: A self-attention based architecture that abandons recurrence in favor of parallelized computations.
  • Informer[zhouInformerEfficientTransformer2021]: An efficient variant of the Transformer specifically designed for long sequence forecasting.
  • Mamba[guMambaLinearTimeSequence2024]: A SSM that exploits linear time-invariant dynamics to efficiently model sequential data.
@@
All models are trained using a batch size of 64. The fixed learning rate is chosen as part of the hyperparameter tuning. The optimization is carried out using the Adam optimizer [DBLP:journals/corr/KingmaB14].
@@
All training is executed on a Kubernetes cluster comprising three nodes, each equipped with 128 CPU cores and 500GB of memory, alongside a total of 4 NVIDIA A30 GPUs. Similar to the approach presented in [steudeEndtoendMLOpsIntegration2024a], we orchestrated the execution of the experiments using Kubeflow pipelines.
@@ 6.1 Baseline Forecasting Performance
Fully Connected DLinear [zengAreTransformersEffective2023] 0.2174 $\pm$ 0.1393 0.2587 $\pm$ 0.1809
Fully Connected MLP [rumelhartLearningRepresentationsBackpropagating1986] 0.2284 $\pm$ 0.1523 0.3916 $\pm$ 0.2460
Convolution TCN [leaTemporalConvolutionalNetworks2017] 0.2470 $\pm$ 0.1550 0.4362 $\pm$ 0.2834
Recurrent LSTM [hochreiterLongShortTermMemory1997] 0.2575 $\pm$ 0.1758 0.5016 $\pm$ 0.3676
Recurrent GRU [choLearningPhraseRepresentations2014] 0.2393 $\pm$ 0.1679 0.4620 $\pm$ 0.3209
Modular RIMs [goyalRecurrentIndependentMechanisms2021] 0.2383 $\pm$ 0.1655 0.3868 $\pm$ 0.2250
Attention Transf. [vaswaniAttentionAllYou2017] 0.2200 $\pm$ 0.1575 0.3576 $\pm$ 0.2123
Attention Informer [zhouInformerEfficientTransformer2021] 0.2203 $\pm$ 0.1550 0.3552 $\pm$ 0.2006
SSM Mamba [guMambaLinearTimeSequence2024] 0.3047 $\pm$ 0.1945 0.7177 $\pm$ 0.7299
@@ 6.2 Robustness Evaluation
Fully Connected DLinear [zengAreTransformersEffective2023] 0.2754 $\pm$ 0.2455
Fully Connected MLP [rumelhartLearningRepresentationsBackpropagating1986] 0.4577 $\pm$ 0.2494
Convolution TCN [leaTemporalConvolutionalNetworks2017] 0.4498 $\pm$ 0.2814
Recurrent LSTM [hochreiterLongShortTermMemory1997] 0.6411 $\pm$ 0.3412
Recurrent GRU [choLearningPhraseRepresentations2014] 0.5948 $\pm$ 0.3543
Modular RIMs [goyalRecurrentIndependentMechanisms2021] 0.5886 $\pm$ 0.2856
Attention Transf. [vaswaniAttentionAllYou2017] 0.4745 $\pm$ 0.2577
Attention Informer [zhouInformerEfficientTransformer2021] 0.4748 $\pm$ 0.2670
SSM Mamba [guMambaLinearTimeSequence2024] 0.5186 $\pm$ 0.3804
@@ 7 Discussion
In this paper, we have introduced a practical definition of robustness tailored specifically to industrial CPS time series and evaluated this definition through a comprehensive forecasting benchmark involving multiple DL architectures. One insight is the strong performance of relatively simple models such as DLinear, confirming earlier findings [zengAreTransformersEffective2023]. However, despite its excellent predictive accuracy on both validation and test datasets, DLinear was highly vulnerable to disturbances, achieving the lowest robustness score among all tested architectures. This shows that standard evaluation methods might not detect overfitting and fail to quantify robustness, which is further affirmed by empirical research in the deployment of machine learning models [paleyesChallengesDeployingMachine2022].